OpenStudy (anonymous):

Using the identity cosec²A = 1 + cot²A, prove that (cosA-sinA+1)/(cosA+sinA-1) = cosecA + cotA

OpenStudy (anonymous):

\[\frac{ \frac{cosa - sina +1 }{sina}}{\frac{cosa + sina -1}{sina}}\]

OpenStudy (anonymous):

This Gives you \[\frac{cota -1 + coseca}{cota +1 - coseca }\]

OpenStudy (anonymous):

Then \[ 1 = cot^2a - cosec^2a \]

OpenStudy (anonymous):

after this...??

OpenStudy (anonymous):

Sorry ...Typo It's \[ 1 = cosec^2a - cot^2a \] So the Whole Equation Becomes \[\frac{(cota + coseca ) ( 1 +cota - coseca)}{1 + cota - coseca }\]

OpenStudy (anonymous):

Then You Get \[cota + coseca \]

OpenStudy (anonymous):

What I did was to replace 1 in Numerator ... through the Identity

OpenStudy (anonymous):

Harkirat Was it Helpful ?

OpenStudy (anonymous):

next???

OpenStudy (anonymous):

Well take \[cota + coseca \]Common and You will see Denominator cancels out .. Giving you the answer

OpenStudy (anonymous):

sorry, my internet went funny..... thanks ishaan for yr help........ medal for u/.......

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Can u help me with one more...?

OpenStudy (anonymous):

Yea Sure :D Post it

OpenStudy (anonymous):

i have proved this by some manipulation but could not proceed from LHS to RHS..... Prove that 1 1 1 1 ---------- - ------ = ------ - ----------- secA + tanA cosA cosA secA - tanA

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

\[\frac{1}{seca - tana } + \frac {1}{tana + seca} = \frac{1}{cosa} + \frac{1}{cosa}\]

OpenStudy (anonymous):

yes, i have used this manipulation to solve it.......... see if you can proceed from LHS and reach RHS without this manipulation......

OpenStudy (anonymous):

So I should Solve it without manipulation ! agh No

OpenStudy (anonymous):

Ok I'm ON it

OpenStudy (anonymous):

I guess Your Internet went Funny Again :D lol

OpenStudy (anonymous):

yeah, it is acting funny today.......

OpenStudy (anonymous):

Hell No....can't do it without manipulation ! NNNoooo lol :D sorry it becomes something too complicated to operate

OpenStudy (anonymous):

ok, thanks......☺ I'll try myself later, right now I am very tired and don't want to tax my brain.....lol !!!